OpenStudy (anonymous):

PLEASE HELP! What is a good website to practice math problems including Finding the equation of a tangent line given the standard equation of a circle and a tangent point on the circle

OpenStudy (anonymous):

your book? idk let me see if my book has any... i dont believe ours went as far as finding it on a circle

OpenStudy (anonymous):

i accidently left my book in my locker, and need to study for a test, so i want some problems to see if i actually can apply what i think i know now

OpenStudy (anonymous):

my book has nothing of the sort it just has lines and stuff =/

OpenStudy (anonymous):

i mean it is basically the same thing though

OpenStudy (anonymous):

can you post one of the ones your book has and let me see how close it is to what im trying to do, please?

OpenStudy (anonymous):

well a circle is a little bit different as you'd have to solve for y

OpenStudy (anonymous):

Find an equation of the tangent line to the graph of f at the given point \[f(x)=x+\frac{4}{x}\] at point (4,5)

OpenStudy (anonymous):

basically the only difference between this question and the circle is a circle will come out as like (x-2)^2+(y-5)^2=25

OpenStudy (anonymous):

the thing a circle does is it makes you figure out the slope of the circles radius then use the opposite recipricle of that, makes you put it into point slope form, to find the slope of the tangent point, so it can only be tangent at one spot of the circle.

OpenStudy (anonymous):

what? i'd believe the only thing you'd need to solve a circle is the equation solved for y , the known quadrant in which the point is located, and the dy/dx of the equation

OpenStudy (anonymous):

because if you know where that point is located you know which part of the squareroot you must use

OpenStudy (anonymous):

this is one of the problems i have to answer already as part of my review, so my review says Find the equation of a line tangent to (x+5)^2 + (y-53)^2 =26 (for the equation of a circle) at the point (-10,52) (as the point of tangency) then you go through all of the steps an get y=-5x+2 as the answer

OpenStudy (anonymous):

well he might have you solve it another way but this is what i was thinking... when you get it solved by y itself you should get an answer like